Mark places one - page, four- color advertisements for a new kind of projector in two magazines: Adventure & Cautionary Tales (ACT), which reaches 52,000 people, and Horrors of the Silver Screen (HSS), which delivers a readership of 64,000 per issue. He estimates that 18% of the readers of Horror of the Silver Screen also read Adventures & Cautionary Tales. Assume that only one impression shall be made for one readership.

a) If ACT and HSS charge $1,200 and $1,500 per ad, respectively, what is the CPM for a campaign that puts one ad in each magazine?
Answer #1

Reach of ACT = 52000

Reach of HSS = 64000

Common readers = 18% of HSS readers = 18%*64000 = 11520

Total reach = 52000 + 64000 -11520 = 104480

Total costs for both magazines = 1200 + 1500 = $2700

CPM = (Total cost / Total impressios) * 1000 = 2700/ 104480*1000 = $25.84